Find the measure of an angle whose supplement measures 38 degrees less than three times its complement.
Answer by checkley71(8403)

2 SUPPLEMENTARY=180 DEGREES.
X+(3X-38)=180
4X-38=180
4X=180+38
4X=218
X=218/4
X=54.5 THE SMALLER SUPPLEMENTARY ANGLE.
3*54.5-38
163.5-38
125.5 THE LARGER SUPPLEMENTARY ANGLE.
PROOF
54.5+125.5=180
180=180